What will take the most energy?

(Constants: Heat of fusion of water: 334 J/g

Heat of vaporization of water: 2260 J/g

Specific heat capacity of copper: 0.385 J/g*C

Specific heat capacity of gold: 0.129 J/g*C

Specific heat capacity of water: 4.18 J/g*C)

a.

Increase the temperature of 10 grams of solid gold by 15 degrees

b.

Increase the temperature of 10 grams of solid water (ice) by 12 degrees

c.

Evaporate 10 grams of water

d.

Increase the temperature of 10 grams of copper by 15 degrees.

e.

Freeze 10 grams of water.

Which statement about light is true?

a.

It oscillates back and forth between wave and particle-like behavior.

b.

It exhibits both wave and particle-like behavior at the same time.

c.

It has neither wave nor particle-like behavior.

d.

It behaves as a particle only.

e.

It behaves as a wave only.

According to the Aufbau principle, the orbitals that electrons will occupy first will be

a. those in the p subshell

b. The ones with electron density farthest from the nucleus

c. the highest energy orbitals

d. those in the d subshell

e. the lowest energy orbitals

True or False: According to Arrhenius's theory, all acids increase the hydrogen ion concentration in water.
Answer #1

1) The heats involved in each case are calculated:

a) q = m * cs * ΔT = 10 * 0.129 * 15 = 19.35 J

b) q = 10 * 2.09 * 12 = 250.8 J

c) q = 2260 * 10 = 22600 J

d) q = 10 * 0.385 * 15 = 57.75 J

e) q = 334 * 10 = 3340 J

Option C.

2) Option B. The light exhibits a wave-particle behavior.

3) Option c. The orbital of greater energy is filled first.

4) True. Arrhenius acids grant H + ions.

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.